OpenStudy (shrutipande9):

u=x^y show that d^3u/dx^2dy=d^3u/dxdydx

OpenStudy (shrutipande9):

all the d's replaced by daba...i dnt know how 2 insert that

hartnn (hartnn):

you need to show or need to verify it for u=x^y ?

OpenStudy (shrutipande9):

u=x^y given...i have 2 prove that lhs=rhs in d next eq

hartnn (hartnn):

cool so you know how to find, \(\large \dfrac{\partial u}{\partial y}\) first ?

OpenStudy (shrutipande9):

yep

hartnn (hartnn):

what will be the answer to that ?

OpenStudy (shrutipande9):

how d hell do u insert these signs...:\

OpenStudy (shrutipande9):

x^ylogx

hartnn (hartnn):

i wrote this code inside \ [...\ ] without spaces \large \dfrac{\partial u}{\partial y} and thats correct, so now can you find \(\large \dfrac{\partial [x^y \log x]}{\partial x}\) ?

OpenStudy (shrutipande9):

i guess...i m not sure weder dis is correct..yx^(y-1)logx+x^(y-1)

hartnn (hartnn):

actually, \(\large \large \dfrac{\partial^3 u}{\partial^2x\partial y}=\large \dfrac{\partial }{\partial x} (\dfrac{\partial }{\partial x}[\large \dfrac{\partial u}{\partial y}]))\)

hartnn (hartnn):

does that^ make sense ?

OpenStudy (shrutipande9):

ahhann...this is wat i wanted 2 verify...cool:D

hartnn (hartnn):

yep, and i think now you can also work out the right side..

OpenStudy (shrutipande9):

yep...thanks a ton:) now i know y ppl write so gud testimonials for u:P

hartnn (hartnn):

haha, aadat se majboor :P

OpenStudy (shrutipande9):

abhi chaane ke jhaad pe maat chadh:P but still thx...the thnx is from my sis...it was her problem

hartnn (hartnn):

ohhh...tabhi mai sochu! ye biotech waale partial kab se seekhne lag agye :O

OpenStudy (shrutipande9):

dude..i had all this partial ka stuff in my first yr..isiliye mujhe thoda yaad aa raha tha..juz wanted 2 confirm..

hartnn (hartnn):

aree, ok baba, bhadakna mat plz :P i was just kidding ;)

OpenStudy (shrutipande9):

dnt wrry...i dnt get angry so quickly:)...i m happy...abhi bhi mujhe ye sab yaad aa raha hai..:D

hartnn (hartnn):

cool :)

OpenStudy (shrutipande9):

:)

OpenStudy (shrutipande9):

u gave me medal?? lol:P

hartnn (hartnn):

i did, for making an effort to solve the problem by yourself too, thats the sign of good learner

OpenStudy (shrutipande9):

hahaha...:P thnx neway

hartnn (hartnn):

you're always welcome ^_^

OpenStudy (shrutipande9):

aise aur 2 ques hai....tell me weder 2 go in d same manner or not..plz

hartnn (hartnn):

if same type , then same manner.

OpenStudy (shrutipande9):

ohh ok...

hartnn (hartnn):

and if you get stuck, i'm always here.

OpenStudy (shrutipande9):

:) thnx a ton